Session: June 1996 (session 2)
Frequency: 6 cm
Obs time (UT): 169 EU 16:30(17/06)-03:30(18/06)
Obs time (UT): 169 US 16:30(17/06)-03:30(18/06)
EU Mk3 Tapes/St.: 1
To give feedback on this experiment, click here
Noto (Corrado Trigilio at Tue Jun 18 09:25:03 EDT 1996) Success
No problems reported
Jodrell (David Henstock at Fri Jun 21 09:15:59 EDT 1996) Minor failures
Sorry, but there was no signal in BBC4 throughout this experiment GM27.
Other problems were:
1) From about 23:00-03:00 UT there were problems with the telescope
accepting source commands due to a faulty High Speed Link card. Some
scans in this time period may have been lost as the telescope was not on the right
2) At 02:30 UT it was noticed that formatter clock was out by 1 second.
Clock was then reset. I do not know how much of the previous GM27
data was affected, and to what extent.
Calibration (Fredrik T Rantakyro at Thu Oct 3 11:54:20 EDT 1996) Minor failures
An ANTAB file for this experiment can be found on VLBEER in the